Find the derivatives of each of the following functions:

a. 𝑦 = 5π‘₯2+7π‘₯βˆ’8

b. 𝑓(π‘₯) = 7/π‘₯4

c. 𝑦 = 15/√π‘₯

d. 𝑓(π‘₯) = 2π‘₯7/2βˆ’π‘₯-1/3

e. 𝑦 = (4π‘₯2βˆ’7π‘₯)/π‘₯

f. 𝑓(π‘₯) = 7π‘₯3/√π‘₯



1
Expert's answer
2022-06-03T07:40:17-0400

a.


"y'=(5\ud835\udc65^2+7\ud835\udc65\u22128)'=10x+7"

b.


"y'=(\dfrac{7}{x^4})'=\dfrac{7(-4)}{x^5}=-\dfrac{28}{x^5}"

c.


"y'=(\dfrac{15}{\sqrt{x}})'=\dfrac{15(-1/2)}{x^{3/2}}=-\dfrac{15}{2x^{3/2}}"

d.


"f'(x)=(2𝑥^{7/2}−𝑥^{-1/3})'=2(\dfrac{7}{2})x^{5/2}-(-\dfrac{1}{3})x^{-4/3}"

"=7x^{5/2}+\dfrac{1}{3}x^{-4/3}"

e.


"y'=((4x^2-7x)\/x)'=(4x-7)'=4"



f.


"f'(x)=( 7𝑥^3/√𝑥)'=7(x^{5/2})'=\dfrac{35}{2}x^{3/2}"